# PCGBandit: One-shot acceleration of transient PDE solvers via online-learned preconditioners

### v2

@@ 4 Evaluation
To assess the effectiveness of PCGBandit at accelerating OpenFOAM simulations, we compare it to running PCG with two baseline preconditioners: incomplete Cholesky (DIC) and multigrid (GAMG). For the latter we select the most commonly used smoother in the tutorials (DIC+Gauss-Seidel) as the default, [^7] The other GAMG options we tune have defaults provided by OpenFOAM. and for fair comparison with the multigrid options in the PCGBandit search space we also implement a backstop and turn off agglomeration caching. In addition to these baselines, when cost permits we also compare to preconditioners drawn uniformly at random from the configuration space described in Section 3.3, as well as to the best-in-hindsight preconditioner from the same.
In the remainder of this section we first detail the main results, including wallclock improvements and the number of timesteps needed for PCGBandit to be useful. Then we investigate the specific configurations found by the method, demonstrate its performance using deterministic cost estimators, and discuss the effect of the number of configurations.
Figure 3: Performance profiles (Dolan and Moré, 2002) comparing PCGBandit to two standard baselines: incomplete Cholesky and geometric-algebraic multigrid with the defaults described in Section 4. The curves report the fraction of the six evaluation tasks from Table 1 for which the corresponding method is $\tau$-suboptimal, i.e. has $\tau\times$ greater wallclock cost, relative to the best method of the three on each task. The PCGBandit curve being in the upper left implies it is optimal or near-optimal on all evaluated tasks.
Figure 4: Break-even timesteps for each simulation setting and every seed. These timesteps are defined to be the one after which PCGBandit’s cumulative cost is smaller than that of both baselines (DIC and GAMG) for the rest of the simulation. This indicates the minimum number of timesteps needed for PCGBandit to be useful in a specific setting. Apart from the 3D DNS simulation—where DIC has roughly the same performance as the best configuration and so PCGBandit never does better—the results demonstrate that the method usually breaks even around a thousand timesteps, and often needs as little as a hundred.
@@
4.1 Main results
Each simulation is run for ten trials each on 2.8 GHz Intel Cascade Lake CPUs. The trials measure the total wallclock, the subset of that time spent on solving SPD systems using PCG, and the subset of the latter spent on learning the preconditioner distribution (where applicable). We start by reporting aggregate results across all the tasks in Table 1 in Figure 3, where we chart the performance profiles (Dolan and Moré, 2002) of PCGBandit and the two static preconditioner baselines. As $\tau\geq 1$ varies, these curves track the fraction of tasks on which each method is at most $\tau$-suboptimal. The results demonstrate that PCGBandit is optimal or near-optimal on all tasks, with both the multigrid and the incomplete Cholesky baselines being around $1.5\times$ or more slower on three tasks each in terms of total wallclock. This effect is even more pronounced when counting only the time spent running PCG, where PCGBandit can be more than $4\times$ faster than both methods.
We next disaggregate the same results in Figure 5, which depicts the total wallclock as a sum of the PCG cost, the learning cost, and other costs. Here we also have the random sampling baseline in addition to the two static preconditioner baselines, and for the 3D DNS and Pitz-Daily simulations we also show the best configuration from the PCGBandit search space. We see that PCGBandit only fails to beat the two static baselines on the simplest task, 3D DNS, where incomplete Cholesky is near optimal. It nevertheless comes reasonably close to the performance of the best-in-hindsight configuration in both settings where it is known, as expected from the theory. Lastly, note that learning costs are so negligible that they are not discernible on the graphs, except for on the Stefan problem, which is distinct in having many cheap iterations.
Finally, in Figure 4 we plot for each seed of every setting its break-even timestep, defined as the timestep $t\in[T]$ after which the cumulative cost up to $t$ is always lower for PCGBandit than it is for the two baselines. Because our method is anytime—its specification does not depend on the total number of timesteps TT—the breakeven timesteps indicate the number of linear system instances a simulation needs to have before learning is useful. While it varies significantly by setting, usually one thousand steps is enough data for PCGBandit, and in some cases very few systems are needed. This can provide some indication of what simulations can benefit from PCGBandit, although as explored in Section 4.4 this heavily depends on the number of configurations $d$.
@@ 4.2 A closer look at within-simulation learning
Figure 7: Heatmaps of the probabilities assigned to each configuration at each solve of the momentum equation (left) and the electric potential (right) on the Shercliff flow MHD simulation. The highlighted configuration denotes the one with the largest total probability assigned across solves.
@@
Figure 9: Heatmaps of the probabilities assigned to each configuration at each solve of the momentum equation (left) and the electric potential (right) on the fringing B-field MHD simulation. The highlighted configuration denotes the one with the largest total probability assigned across solves.
We now take a closer look at the learning process on the two MHD simulations, Shercliff flow and fringing B-field; as shown in Figure 5 (right), the improvement due to PCGBandit is much more substantial on the latter. In Figure 6 we look at performance separately across the two main SPD equations of Shercliff flow and see that while PCGBandit noticeably improves the solving of the momentum equation (left), it performs worse than the baselines on electric potential (right). Much of this is due to large costs incurred on systems very early in the simulation, while later in the simulation costs are similar to those of the baselines. From Figure 7 we see that the learner is more uncertain about the optimal configuration when solving for potential, although it does converge to a choice similar to the better of the two baselines (DIC rather than multigrid). Notably, Shercliff flow was one of the worst simulations for random sampling when comparing it to a fixed baseline (see Figure 5, top right), making it not surprising that a method such as PCGBandit that begins with a uniform distribution may struggle more there. To fix this issue of trying many related and bad GAMG configurations, future work can consider better initial distributions or bandit algorithms that take advantage of known relationships between methods (Valko et al., 2014), e.g. to use the fact that different multigrid configurations will perform more similar to each than to different incomplete Cholesky variants.
In contrast to the probability evolution when simulating Shercliff flow, Figure 9 shows that for the fringing B-field simulation PCGBandit’s distributions for both equations are rapidly dominated by single configurations. While Figure 8 shows no overall improvement over DIC when solving for electric potential, in this simulation the compute is dominated by the momentum solve (c.f. Table 1), which is significantly accelerated.
@@ 4.3 Reproducible learning via deterministic cost estimates
Figure 10: Performance profiles (Dolan and Moré, 2002) comparing regular PCGBandit and deterministic PCGBandit, where the latter uses Equation 4 to estimate the cost of each solve. We evaluate this in the setting of choosing between ICTC thresholds, and so the baselines are DIC (incomplete Cholesky) and sampling random configurations. The results demonstrate that using deterministic cost estimators is nearly as good as using the wallclock time.
Figure 11: We consider using PCGBandit to tune the ICTC threshold, with the configuration space thus containing only DIC and ICTC with one to thirty-two different thresholds. The first five plots depict the performance of PCGBandit as a function of the number of thresholds (i.e. configurations) considered; here we do not consider the fringing B-field simulation due to cost. The last plot depicts the break-even timestep for each simulation, defined here as the timestep after which the average across random seeds of the cumulative cost of PCGBandit is smaller than that of the DIC baseline.
As noted in Section 3.4, using the wallclock cost of each solve as the bandit feedback leads to runs that are unreproducible even when fixing the random seed, which makes debugging difficult. However, we can sometimes get a strong-enough signal from the number of PCG iterations taken by doing a rough analysis of the floating pointing operations (FLOPs) used by the selected preconditioner at each step. To test the efficacy of this, we run PCGBandit over a restricted search space of eight configurations—ICTC with one of eight different thresholds or DIC—and use the following estimate for the cost at time $t$ given the number of iterations $k_{t}$ that were required to solve $\mathbf{A}_{t}\mathbf{x}=\mathbf{b}_{t}$:
@@ 4.4 The effect of the number of configurations
As discussed in Section 3.2, the number of configurations $d$ considered by PCGBandit has a significant effect on the worst-case regret. We now investigate how this manifests in practical simulations by using PCGBandit to tune only the ICTC threshold while varying the number of thresholds under consideration (we also keep the usual incomplete Cholesky preconditioner DIC as an option). Specifically, we study the effect of learning across $2^{p}$ different thresholds for integers $p\in\{0,\dots,5\}$. Figure 11 demonstrates that the number of thresholds considered has in some cases a dramatic effect on the performance, but the best performance is typically obtained at five or nine configurations (i.e. four or eight thresholds plus DIC). As we might expect from the worst-case analysis, the bottom right plot in Figure 11 also shows that on the harder (for PCGBandit) simulations the break-even point is typically increasing as a function of the number of configurations $d$.

### v3

@@ 4 Evaluation
To assess the effectiveness of PCGBandit at accelerating OpenFOAM simulations, we compare it to running PCG with two baseline preconditioners: IC(0) and multigrid. For the latter we select the smoother used most often in the tutorials (IC(0)+Gauss-Seidel) as the default, [^8] All other multigrid options have defaults provided by OpenFOAM. and for fair comparison with the options in the PCGBandit search space we also implement a backstop and turn off agglomeration caching. In addition to these baselines, we compare PCGBandit to two alternative bandit methods: (1) Tsallis-INF with importance-weighted (IW) cost estimation, which was used by Khodak et al. (2024), and (2) Thompson sampling, specifically its Gaussian variant (Agrawal and Goyal, 2017). When cost permits we also compare to preconditioners drawn uniformly at random from the configuration space described in Section 3.3, as well as to the best-in-hindsight preconditioner from the same.
We next detail the main results, including wallclock improvements and the number of timesteps needed for PCGBandit to be useful. Then we investigate the specific configurations found by the method, demonstrate its performance using deterministic cost estimators, and discuss the effect of search space size.
Figure 4: Performance profiles (Dolan and Moré, 2002) comparing bandit algorithms to two standard baselines: incomplete Cholesky and geometric-algebraic multigrid with the defaults described in Section 4. The curves report the fraction of the six evaluation tasks from Table 1 for which the corresponding method is $\tau$-suboptimal, i.e. has $\tau\times$ greater wallclock cost, relative to the best method of the three on each task. The PCGBandit (Tsallis-INF (RV)) curve being in the upper left implies it is optimal or near-optimal on all evaluated tasks.
4.1 Main results
Each simulation is run for ten trials each on 2.8 GHz Intel Cascade Lake CPUs. The trials measure the total wallclock, the subset of that time spent on solving SPD systems using PCG, and the subset of the latter spent on learning (where applicable). We start by reporting aggregate results across tasks in Table 1 in Figure 4, where we chart the performance profiles (Dolan and Moré, 2002) of the two static preconditioner baselines and three bandit methods. As $\tau\geq 1$ varies, these curves track the fraction of tasks on which each method is at most $\tau$-suboptimal. The results demonstrate that PCGBandit (Tsallis-INF (RV)) is optimal or near-optimal on all tasks, with both multigrid and IC(0) being around $1.5\times$ or more slower on three tasks each in terms of total wallclock. This effect is even more pronounced when counting only the time spent running PCG, where PCGBandit can be more than $4\times$ faster than both. Lastly, among our choices of learning algorithms, Tsallis-INF (IW) is only slightly worse than PCGBandit, while Thompson sampling underperforms more significantly, perhaps due to its stationarity assumption on the costs.
@@ 4.1 Main results
We next disaggregate the same results in Figure 5, which depicts the total wallclock as a sum of the PCG cost, the learning cost, and other costs. Here we add the random baseline to the two static preconditioner baselines, and for the 3D DNS and Pitz-Daily simulations we show the best configuration from the PCGBandit search space. We see that PCGBandit only fails to beat the two static baselines on the simplest task, 3D DNS, where IC(0) is near optimal. It nevertheless comes reasonably close to the performance of the best-in-hindsight in both settings where it is known, as expected from the theory. Lastly, note that learning costs are so negligible that they are not discernible on the graphs.
Finally, in Figure 6 we plot for each bandit algorithm the break-even timestep of each trial, defined as the timestep $t\in[T]$ after which the cumulative cost is always lower for the learner than it is for the two baselines. Because the methods are anytime—they do not take $T$ as an input parameter—the breakeven timesteps indicate the number of linear system instances a simulation needs to see before learning is useful. For PCGBandit (Tsallis-INF (RV)), one thousand steps is usually enough data, although in some cases far fewer systems are needed. Tsallis-INF (IW) and Thompson sampling typically need more, except the latter outperforms on the turbulence simulation (3D DNS); this is likely explained by the stationarity of that setting’s costs (see Figure 3). Overall, the chart provides some indication of what kind of simulations can benefit from learning, although as explored in Section 4.4 the break-even timestep can heavily depend on the number of configurations $d$.
Figure 6: Break-even timesteps for each simulation setting, seed, and bandit algorithm. These timesteps are defined to be the one after which the learner’s cumulative cost is smaller than that of both baselines (IC(0) and multigrid) for the rest of the simulation. This indicates the minimum number of timesteps needed to be useful in a specific setting. The results demonstrate that PCGBandit (Tsallis-INF (RV)) usually breaks even around a thousand timesteps, and often needs a hundred or less.
@@ 4.2 A closer look at within-simulation learning
We now take a closer look at the learning process on the two MHD simulations, Shercliff flow and fringing B-field; as shown in Figure 5, the improvement due to PCGBandit is much more substantial on the latter. In Figure 7 we look at performance separately across the two main Shercliff flow equations and see that while PCGBandit noticeably improves the solving of the momentum equation (left), it performs poorly on electric potential (right). Much of this is due to large costs incurred very early in the simulation, while later PCGBandit’s costs are similar to those of the baselines. As shown in Figure 3, early solves dominate the runtime on Sherclif flow, making it a hard case for a bandit algorithm since it must do well without having seen much data. Notably, Shercliff flow was also one of the worst simulations for random sampling when comparing it to a fixed baseline (see Figure 5, top right), making it not surprising that a method such as PCGBandit that begins with a uniform distribution may struggle more there. To fix this issue of trying many bad configurations, future work can consider better initial distributions or bandit algorithms that take advantage of known relationships between methods (Valko et al., 2014), e.g. to use the fact that different multigrid configurations will perform more similar to each other than to different incomplete Cholesky variants.
@@
Figure 8: Heatmaps of probabilities assigned to each configuration at each solve of the momentum equation (left) and the electric potential (right) on the Shercliff flow simulation. The highlighted configuration is the one with the largest total probability across solves.
@@
Figure 10: Heatmaps of probabilities assigned to each configuration at each solve of the momentum equation (left) and the electric potential (right) on the fringing B-field simulation. The highlighted configuration is the one with the largest total probability across solves.
In the case of the fringing B-field simulation, where PCGBandit performs much better, Figure 10 demonstrates more rapid convergence to a probability distribution dominated by a single configuration than in the case of Shercliff flow (cf. Figure 8). Furthermore, while Figure 9 shows no overall improvement over IC(0) on electric potential, in this simulation the compute is dominated by the momentum solve (c.f. Table 1), which is significantly accelerated.
@@ 4.3 Reproducible learning via deterministic cost estimates
Figure 11: Performance profiles (Dolan and Moré, 2002) comparing regular PCGBandit and deterministic PCGBandit, where the latter uses Equation 4 to estimate the cost of each solve. We evaluate this in the setting of choosing between ICT thresholds, and so the baselines are IC(0) and sampling random configurations. The results demonstrate that using deterministic cost estimators is nearly as good as using the wallclock time.
Figure 12: The break-even timestep for each simulation, defined here as the timestep after which the median across random seeds of the cumulative cost of PCGBandit is smaller than that of the IC(0) baseline. As predicted by the $\mathcal{O}(\sqrt{dT})$ regret, the number of timesteps required to break even typically increases with the number of configurations.
Figure 13: We consider using PCGBandit to tune the ICT threshold, with the configuration space thus containing only IC(0) and ICT with one to thirty-two different thresholds. Here we plot the performance of PCGBandit as a function of the number of thresholds (i.e. configurations) considered.
As noted in Section 3.4, using the wallclock cost of each solve as the bandit feedback leads to runs that are unreproducible even when fixing the random seed, which makes debugging difficult. However, we can sometimes get a strong-enough signal from the number of PCG iterations taken by doing a rough analysis of the floating pointing operations (FLOPs) used by the selected preconditioner at each step. To test the efficacy of this, we run PCGBandit over a restricted search space of eight configurations—ICT with one of eight different thresholds or IC(0)—and use the following estimate for the cost at time $t$ given the number of iterations $k_{t}$ that were required to solve $\mathbf{A}_{t}\mathbf{x}=\mathbf{b}_{t}$:
@@ 4.4 The effect of the number of configurations
As discussed in Section 3.2, the number of configurations $d$ considered by PCGBandit has a significant effect on the worst-case regret. We now investigate how this manifests in practical simulations by using PCGBandit to tune only the ICT threshold while varying the number of thresholds under consideration (we also keep the usual incomplete Cholesky preconditioner IC(0) as an option). Specifically, we study the effect of learning across $2^{p}$ different thresholds for integers $p\in\{0,\dots,5\}$. Figure 13 demonstrates that the number of thresholds considered has in some cases a dramatic effect on the performance, but the best performance is typically obtained at five or nine configurations (i.e. four or eight thresholds plus IC(0)). As we might expect from the worst-case analysis, the bottom right plot in Figure 13 also shows that on the harder (for PCGBandit) simulations the break-even point is typically increasing as a function of the number of configurations $d$.